What are the mass and weight of (a) a 1420-1b snowmobile and (b) a 412=kg heat pump?

Grade:11

1 Answers

Aditi Chauhan
8 years ago

We have to find out the mass and weight for both the parts.
(a) Weight of snow mobile, W = 1420-lb
= (1420-lb) (4.448N/lb)
= 6316.16 N
Rounding off to three significant figures, weight of snow mobile would be, 6320 N.
Weight of a body is equal to the mass of the body times (m) acceleration due to gravity (g) on that surface.
W = mg …… (1)
We know that that, the value of acceleration due to gravity (g) on the surface of earth is 9.81 m/s2.
So, from equation (1), m = W /g
= (6320 N/9.81m/s2)
= (6320 N/9.81m/s2) (1kg.m/s2 /1 N)
= 644.24 kg
Rounding off to three significant figures, mass of snow mobile would be, 644 kg.
Thus from the above observation we conclude that, mass of snow mobile is 644 kg and weight of snow mobile is 6320 N.
(b) Mass of heat pump, m = 412-kgWeight of a body is equal to the mass of the body times (m) acceleration due to gravity (g) on that surface.
W = mg …… (2)
We know that that, the value of acceleration due to gravity (g) on the surface of earth is 9.81 m/s2.
So putting the value of m and g in equation (2), weight of the heat pump would be,
W = mg
= (412-kg) (9.81 m/s2)
= (4041.72 kg. m/s2) (1 N/1 kg. m/s2)
= 4041.72 N
Rounding off to three significant figures, mass of heat pump would be, 4040 kg.
Thus from the above observation we conclude that, mass of heat pump is 412-kg and weight of snow mobile is 4040 N.